OpenStudy (anonymous):

if f(x) = sqrt (x-1) has an inverse function, use the inverse to find (f^-1) (2)

OpenStudy (anonymous):

y= sqrt(x-1) square both sides y^2 = x-1 x = y^2 - 1 so : f-1(x) = x^2 + 1 f-1(x) = 2^2 +1 = 5

OpenStudy (anonymous):

if you want to check yourself then you plug y(5) and should get 2 ..

OpenStudy (anonymous):

yep thanks man!